OpenStudy (anonymous):

I'm a little confused what I'm supposed to do after this point: 12a^-1b^-1. Do I have to invert them or whatever?

jimthompson5910 (jim_thompson5910):

Yes, a^-1 = 1/a and b^-1 = 1/b

jimthompson5910 (jim_thompson5910):

So 12a^-1b^-1 = 12(1/a)(1/b) = 12/(ab)

jimthompson5910 (jim_thompson5910):

Oh sry, skipped a step a^-1 = 1/a^1 = 1/a, but hopefully you get the idea

OpenStudy (anonymous):

Great. Thank you c:

OpenStudy (anonymous):

\[a^{-b} = (\frac{1}{a})^b\]

jimthompson5910 (jim_thompson5910):

you're welcome